# Non-Invasive Medical Digital Twins using Physics-Informed Self-Supervised Learning

### v1

@@ Abstract
A digital twin is a virtual replica of a real-world physical phenomena that uses mathematical modeling to characterize and simulate its defining features. By constructing digital twins for disease processes, we can perform in-silico simulations that mimic patients’ health conditions and counterfactual outcomes under hypothetical interventions in a virtual setting. This eliminates the need for invasive procedures or uncertain treatment decisions. In this paper, we propose a method to identify digital twin model parameters using only noninvasive patient health data. We approach the digital twin modeling as a composite inverse problem, and observe that its structure resembles pretraining and finetuning in self-supervised learning (SSL). Leveraging this, we introduce a physics-informed SSL algorithm that initially pretrains a neural network on the pretext task of solving the physical model equations. Subsequently, the model is trained to reconstruct low-dimensional health measurements from noninvasive modalities while being constrained by the physical equations learned in pretraining. We apply our method to identify digital twins of cardiac hemodynamics using noninvasive echocardiogram videos, and demonstrate its utility in unsupervised disease detection and in-silico clinical trials.
Machine Learning, ICML
@@ 1 Introduction
With increasing health data availability, there is excitement about refining physics-based models of human body systems to be patient specific, as personalized physical models can provide the basis for in-silico experiments and timely diagnosis (Gray & Pathmanathan, 2018; Ienca et al., 2018; Pappalardo et al., 2019a; Gottesman et al., 2023). This personalized vision of healthcare has given rise to virtual patients constructed with data-tuned models, known as digital twins. Originally an engineering concept, digital twins have recently been realized as a resource in healthcare (Bruynseels et al., 2018; Björnsson et al., 2020). The concept combines data-driven approaches with mechanistic or simulation techniques, serving as a bidirectional map between health and intervention. Medical digital twins have been applied and developed broadly in applications ranging from cellular mechanics (Filippo et al., 2020) to the development of whole body and human digital twins (Okegbile et al., 2023; Wang et al., 2024), with aims ranging from tailoring cardiovascular interventions (Hirschvogel et al., 2019) to agent-based trauma care management systems (Croatti et al., 2020). Both academic and industry efforts are prolific with digital twins having been developed for various medical problems (Canzoneri et al., 2021; Subramanian, 2020; Voigt et al., 2021; Corral-Acero et al., 2020; Erol et al., 2020; Sun et al., 2023).
Figure 1: Digital twins for the cardiac hemodynamics. Left: Representation of the cardiovascular system with its key components. Middle: Our proposed physics-informed self-supervised learning approach for learning digital twins from measurements of the heart. Right: Examples of digital twin models of cardiac hemodynamics based on hydraulic or electric representations.
Figure 1: Digital twins for the cardiac hemodynamics. Left: Representation of the cardiovascular system with its key components. Middle: Our proposed physics-informed self-supervised learning approach for learning digital twins from measurements of the heart. Right: Examples of digital twin models of cardiac hemodynamics based on hydraulic or electric representations.
In this paper, we propose a methodology for training digital twins with medical image data directly. Medical imaging, as a non-invasive modality, plays an important role in describing patient health. The inductive task of determining patient health status from images is a core medical challenge with growing computational interest and rapidly developing methodology (Castiglioni et al., 2021; Razzak et al., 2018; Ouyang et al., 2020). In the era of personalized medicine, medical imaging informs complex intervention procedures and mathematical models. Examples include the application of digital twins in oncology (Wu et al., 2022). Medical imaging is used to calibrate physics-based models using patient specific rates or model parameters extracted from the obtained images (Baldock et al., 2013; Scheufele et al., 2021; Le et al., 2016). The central role of medical imaging in screening, diagnosing, staging, guiding therapy, and monitoring also suggests it as an important, direct source of patient-specific learnable measures. Medical imaging provides serial observations of the spatiotemporal variations in the physical state of the patient, which could itself be used directly as input data in models developing digital twins. (Ouyang et al., 2020) demonstrate the ability to predict left ventricular ejection fraction, which is typically hand-labeled by cardiologists, directly from ultrasound video data using convolutional neural networks. The direct use of medical imaging for developing digital twin technology holds great potential for non-invasively estimating patient measurements that traditionally require invasive procedures.
There has also been a large amount of recent work developing methods to learn parameters and solutions of mathematical models leveraging neural network approaches in physics-informed neural networks (PINNs) (Raissi et al., 2019), in addition to other methods of bringing physical assumptions into machine learning for broad applications including hemodynamics and flow problems (von Rueden et al., 2023; Cuomo et al., 2022). We propose that combining these efforts with image prediction tasks directly can provide the basis for non-invasively predicting patient specific physics-based models and the creation of digital twins for precision medicine (Rasheed et al., 2020). Digital twins that combine medical imaging and physical models have the potential to improve upon methods to non-invasively estimate broad categories of patient measurements that traditionally require invasive or expensive procedures for 1) evaluating patient health status and prognosis, 2) providing timely and accurate diagnoses, and 3) performing in-silico trials for determining treatments and guiding interventions.
@@
We study obtaining latent parameters of a physics-based model from image measurements directly as an inverse problem broken into composite tasks. Learning patient specific parameters for a physics-based model from a complete description of patient physical states and a known forward model is an inverse problem. This task can be modeled using, for example, a PINN approach (van Herten et al., 2022; Kissas et al., 2020; Herrero Martin et al., 2022). The process of obtaining patient physical states from non-invasive medical imaging is a second inverse problem with an unknown forward model. Obtaining patient specific latent parameters from non-invasive medical imaging directly combines these two inverse problems and could be modeled by learning the unknown process components. Large numbers of labeled data pairs of images and parameters for the full inverse task would need to be obtained from invasive procedures to use supervised learning. Instead, we leverage labeled data pairs of images and non-invasively measurable patient physical states. We first train a neural network using supervised learning with synthetic data to approximate the forward physics-based model as a pre-training task. We then used this fixed network to fine-tune the learning of a solution to the composite inverse problem from the labeled measurable data pairs. Our methods are then extendable to include exogenous intervention parameters for the development of in-silico experiments.
@@
We apply our methodology in the setting of predicting patient specific cardiac pressure-volume loops from echocardiography. The relationship between left-ventricle pressure and volume is an important description of cardiac function (Kass et al., 1988). Pressure-volume catheter measurements are highly invasive, challenging their routine clinical use (Burkhoff et al., 2005; Abawi et al., 2022; Wei et al., 2014). Cardiac ejection fraction, the ratio between left-ventricular stroke volume and end-diastolic volume, while more routinely measured as a proxy for cardiac function, does not take into account cardiac pressure, preventing diagnosis of many severe cardiac conditions and providing motivation to develop digital twins for pressure-volume relations (Westermann et al., 2008; Marwick, 2018; Potter & Marwick, 2018). Cardiovascular pressure and volume are commonly modeled using hydraulic analogy models, representing the hydraulic system by electronic circuits, characterized by lumped parameters corresponding to system attributes. Systems of differential equations describe the relationships between circuit parameters and volume and pressure states. A model developed in (Simaan et al., 2008) relates left-ventricular pressure and volume with and without the addition of a left-ventricular assistance device (LVAD). We train a neural network to learn to map between parameters and solution states of this model and use this learning to fine-tune a convolutional neural network that learns patient specific model parameters using echocardiography frames.
@@ 2.1 Problem Setup
The model $\mathcal{M}$ is assumed to be known and would typically be defined by a set of differential equations with a physically interpretable parameter set $\theta$. In Section 3, we examine a particular instance of $\mathcal{M}$ in the context of cardiovascular hemodynamics. The level of granularity in simulating the (true) physiological process depends on the fidelity of the physics-based model $\mathcal{M}$.
In addition to modeling the natural state of the physiological process, we also consider physics-based models of clinical interventions (e.g., medical devices). To model an interventional process, we consider an augmented version of the physics-based model ~$\widetilde{\mathcal{M}}:\Theta\times\mathcal{U}\to\mathcal{X}$, which maps the endogenous parameters $\theta\in\Theta$ of the patient and an exogenous intervention $u\in\mathcal{U}$ to a physiological state $\mathbf{x}\in\mathcal{X}$. The augmented model describes how the intervention $u$ alters the physiological states of a patient with parameters $\theta$—in the causal inference literature, this class of models is known as structural causal models (SCMs) (Pearl, 2009). In Section 3, we consider an instance of $\widetilde{\mathcal{M}}$ that models an implantable pump for circulating blood through the body.
To identify the model $\mathcal{M}(\theta)$ and simulate the states $\mathbf{x}$, one would need to conduct a number of invasive procedures for each patient. Here, we assume that we have access to a measurement $\mathbf{y}_{i}$ for each patient $i$, which reflects some aspects of the underlying physiological process. We assume that $\mathbf{y}$ is acquired through a non-invasive procedure that might be conducted routinely in clinical practice (e.g., an electrocardiogram). The relationship between the physiological state $\mathbf{x}$ and the non-invasive measurement $\mathbf{y}$ is given by
@@
Since the physical model $\mathcal{M}$ is described by a known system of differential equations, one can use numerical methods to approximate $\mathcal{M}^{-1}$. However, we do not directly observe the state $\mathbf{x}$, but rather a measurement $\mathbf{y}$ obtained by passing $\mathbf{x}$ through the unknown forward model $\mathcal{K}$. Thus, obtaining the physical parameters from $\mathbf{y}$ entails solving the composite inverse problem given by:
@@
Solving this problem entails two challenges. First, we do not know the forward model $\mathcal{K}$, hence we cannot use numerical methods to approximate its inverse. Second, we cannot directly learn the inverse map $\mathcal{F}^{-1}$ using supervised learning since this requires access to a labeled dataset $\{(\mathbf{y}_{i},\theta_{i})\}_{i}$. Obtaining such a dataset requires conducting a large number of invasive procedures to acquire training data, antithetical to the purpose of modeling digital twins in the first place.
An alternative way to learn the inverse map $\mathcal{F}^{-1}$ is to break down the composite inverse problem in (5) into separate inverse problems, and only learn the inverse of the unknown map $\mathcal{K}^{-1}$ using examples of the form $\{(\mathbf{y}_{i},\mathbf{x}_{i})\}_{i}$. While we do not have direct access to the patient physiological state $\mathbf{x}_{i}$, the noninvasive observation $\mathbf{y}_{i}$ typically measures physiological quantities that can also be derived from the true state $\mathbf{x}_{i}$. We assume that $\bar{\mathbf{x}}$ is a physiological variable that can be derived from both $\mathbf{x}$ and $\mathbf{y}$, i.e.,
@@
Given the setup above, our digital twin modeling problem can be formulated as follows. Given a dataset of noninvasive measurements for $n$ patients, $\mathcal{D}=\{\mathbf{y}_{i}\}_{i=1}^{n}$, and a physiological variable of interest $\bar{\mathbf{x}}_{i}=m(\mathbf{x}_{i})=g(\mathbf{y}_{i})$, our goal is to train a model that can identify the underlying physics-based twin $\mathcal{M}(\theta_{n+1})$ for a new patient $n+1$ based solely on their corresponding noninvasive measurement $\mathbf{y}_{n+1}$.
@@ 2.2 Physics-Informed Self-Supervised Learning
The motivation for training the model $\widehat{f}$ is usually to automate the collection of physiological variables from medical images in a clinical workflow (Stoyanov et al., 2018). These models follow a fully data-driven approach and do not incorporate our knowledge of how the physiological processes being measured function on the biological or physical levels. One could think of the supervised model $\widehat{f}$ as a data-driven model of the function $g(.)$ in (6).
@@
$\widehat{\phi}_{\mathcal{M}}=\arg\min_{\phi}\frac{1}{\tilde{n}}\sum_{j=1}^{\tilde{n}}\ell(\phi(\tilde{\theta}_{j}),\tilde{\mathbf{x}}_{j}).$ (10)
@@
Step 2: Physics-Guided Finetuning. Given the observed dataset $\mathcal{D}=\{(\mathbf{y}_{i},\bar{\mathbf{x}}_{i})\}_{i}$, we train another neural network to predict the physical parameters $\theta$ from the observed measurements $\mathbf{y}$ using the following loss function
@@
Interpretation of Physics-Informed SSL. Our proposed algorithm structurally resembles the SSL paradigm, where, akin to SSL, we decompose our model into a “backbone” architecture and a task-specific “head” (Misra & Maaten, 2020). However, there are fundamental conceptual differences between standard SSL approaches and ours. In standard SSL, the backbone is a high-capacity model pretrained on a pretext task using unlabeled data to derive a general-purpose representation transferable to many downstream tasks, while the low-capacity head is finetuned for the specific task. In our physics-informed approach, SSL serves to constrain rather than enhance flexibility. We first pretrain the low-capacity head to distill the laws of physics, and then finetune the high-capacity backbone model to learn a mapping that aligns with the frozen head to produce the observed physiological variables $\mathbf{x}$ while respecting the physical laws. In doing so, we force the backbone model to learn a mapping from the measurement $\mathbf{y}$ to the physical parameters $\theta$.
The concept of distilling a physics-based dynamic model into a data-driven one draws from engineering methodologies that develop surrogate models to simplify computational complexity using simulated data (e.g., (Pfrommer et al., 2018)). This data-driven approach mirrors other supervised learning generative strategies for solving inverse problems or system identification with known dynamics in control theory (Ongie et al., 2020; Werbos, 1991). Importantly, physics-informed SSL scales efficiently with increasingly complex and higher fidelity physical models, as its computational requirements involve one offline process (Eq. (9) and (10)) once the synthetic data is acquired.
Figure 2: Illustration of physics-informed self-supervised learning of digital twins for cardiovascular hemodynamics. (a) Pictorial depiction of the two-step algorithm proposed in Section 2.2. (b) Five state lumped-parameter electric circuit model of cardiac hemodynamics from (Simaan et al., 2008). Here $\mathbf{x}(t)=[x_{1}(t),x_{2}(t),x_{3}(t),x_{4}(t),x_{5}(t)]=[P_{LV}(t),P_{LA}(t),P_{A}(t),P_{Ao}(t),Q(t)]$ describe the voltages $x_{1},x_{2},x_{3},x_{4}$ or pressures in the left-ventricle, left atrium, arteries, and aorta, respectively, and total flow $x_{5}$. The LVAD is modeled through an electric circuit connected to the digital twin via a switch. An LVAD intervention is applied if the switch is closed.
Figure 2: Illustration of physics-informed self-supervised learning of digital twins for cardiovascular hemodynamics. (a) Pictorial depiction of the two-step algorithm proposed in Section 2.2. (b) Five state lumped-parameter electric circuit model of cardiac hemodynamics from (Simaan et al., 2008). Here $\mathbf{x}(t)=[x_{1}(t),x_{2}(t),x_{3}(t),x_{4}(t),x_{5}(t)]=[P_{LV}(t),P_{LA}(t),P_{A}(t),P_{Ao}(t),Q(t)]$ describe the voltages $x_{1},x_{2},x_{3},x_{4}$ or pressures in the left-ventricle, left atrium, arteries, and aorta, respectively, and total flow $x_{5}$. The LVAD is modeled through an electric circuit connected to the digital twin via a switch. An LVAD intervention is applied if the switch is closed.

### v2

# Med-Real2Sim: Non-Invasive Medical Digital Twins using Physics-Informed Self-Supervised Learning

@@ Abstract
A digital twin is a virtual replica of a real-world physical phenomena that uses mathematical modeling to characterize and simulate its defining features. By constructing digital twins for disease processes, we can perform in-silico simulations that mimic patients’ health conditions and counterfactual outcomes under hypothetical interventions in a virtual setting. This eliminates the need for invasive procedures or uncertain treatment decisions. In this paper, we propose a method to identify digital twin model parameters using only noninvasive patient health data. We approach the digital twin modeling as a composite inverse problem, and observe that its structure resembles pretraining and finetuning in self-supervised learning (SSL). Leveraging this, we introduce a physics-informed SSL algorithm that initially pretrains a neural network on the pretext task of learning a differentiable simulator of a physiological process. Subsequently, the model is trained to reconstruct physiological measurements from noninvasive modalities while being constrained by the physical equations learned in pretraining. We apply our method to identify digital twins of cardiac hemodynamics using noninvasive echocardiogram videos, and demonstrate its utility in unsupervised disease detection and in-silico clinical trials.
@@ 1 Introduction
With increasing health data availability, there is excitement about refining physics-based models of human body systems to be patient specific, as personalized physical models can provide the basis for in-silico experiments and timely diagnosis (26; 30; 54; 24). This personalized vision of healthcare has given rise to virtual patients constructed with data-tuned models, known as digital twins. Originally an engineering concept, digital twins have recently been realized as a resource in healthcare (6; 4). The concept combines data-driven approaches with mechanistic or simulation techniques, serving as a bidirectional map between real data and simulations. Medical digital twins have been applied and developed broadly in applications ranging from cellular mechanics (23) to the development of whole body and human digital twins (49; 75), with aims ranging from tailoring cardiovascular interventions (29) to agent-based trauma care management systems (14). Both academic and industry efforts are prolific with digital twins having been developed for various medical problems (9; 68; 73; 13; 21; 69).
In this paper, we propose a framework for identifying patient-level digital twins using noninvasive medical images. We focus on scenarios where identifying a patient’s digital twin from noninvasive data allows us to simulate other physiological parameters that typically require invasive measurement, such as through catheterization. Such an approach assumes that there is a mapping from rich noninvasive imaging data to patient-specific mechanistic models of physiology; such a mapping is not known to scientists humans but can be learned from data. This hypothesis is supported by previous experimental work (2; 63; 42). However, the challenge lies in the absence of datasets containing both noninvasive data and the corresponding parameters of the underlying physics-based physiological model, hence the problem cannot be solved using standard supervised learning. To this end, we propose a new class of physics-informed neural networks (PINNs) (59) to incorporate inductive biases informed by mechanistic models of patient physiology into neural network architectures.
@@
Summary of contributions. We study obtaining latent parameters of a physics-based model from image measurements directly as an inverse problem. Learning patient specific parameters for a physics-based model from a complete description of patient physical states and a known forward model is an inverse problem. This task can be modeled using, for example, a PINN approach (70; 38; 28). The process of obtaining patient physical states from non-invasive medical imaging is a second inverse problem with an unknown forward model. Obtaining patient specific latent parameters from non-invasive medical imaging directly combines these two inverse problems and could be modeled by learning the unknown process components. Large numbers of labeled data pairs of
Figure 1: Digital twins for cardiac hemodynamics. Left: Illustration of the cardiovascular system. Right: Digital twin models of cardiac hemodynamics based on hydraulic or electric representations.
images and parameters for the full inverse task would need to be obtained from invasive procedures to use supervised learning. Instead, we propose a training approach that structurally resembles self-supervised learning (SSL), and leverages labeled data pairs of images and non-invasively measurable patient physical states. We first train a neural network using supervised learning with synthetic data to approximate the forward physics-based model as a pre-text task. We then freeze this network to fine-tune the learning of a solution to the composite inverse problem from the labeled measurable data pairs. Our methods are then extendable to include exogenous intervention parameters for the development of in-silico experiments. We call this framework “Med-Real2Sim” since it learns virtual simulators of patient phyisology on the basis of real (noninvasive) data.
@@
We apply our methodology in the setting of predicting patient specific cardiac pressure-volume loops from echocardiography. The relationship between left-ventricle pressure and volume is an important description of cardiac function (35). Pressure-volume catheter measurements are highly invasive, challenging their routine clinical use (8; 1; 76). Cardiac ejection fraction, the ratio between left-ventricular stroke volume and end-diastolic volume, while more routinely measured as a proxy for cardiac function, does not take into account cardiac pressure, preventing diagnosis of many severe cardiac conditions and providing motivation to develop digital twins for pressure-volume relations (79; 44; 58). Cardiovascular pressure and volume are commonly modeled using hydraulic analogy models, representing the hydraulic system by electronic circuits, characterized by lumped parameters corresponding to system attributes. Systems of differential equations describe the relationships between circuit parameters and volume and pressure states. A model developed in (65) relates left-ventricular pressure and volume with and without the addition of a left-ventricular assistance device (LVAD). We train a neural network to learn to map between parameters and solution states of this model and use this learning to fine-tune a convolutional neural network that learns patient specific model parameters using echocardiography frames.
@@ 2.1 Problem Setup
The model $\mathcal{M}$ is assumed to be known and would typically be defined by a system of differential equations written $f_{\theta}(\mathbf{x})=0$ with a physically interpretable parameter set $\theta$. One can use numerical methods to solve $f_{\theta}(\mathbf{x})=0$ for $\mathbf{x}$ and construct the solution $\mathcal{M}(\theta_{i})$ for each patient parameter set $\theta_{i}$ but a closed form solution to define the map $\mathcal{M}$ may not exist even though the model is known, such as in the case where $f_{\theta}=0$ has no closed form solution. The level of granularity in simulating the (true) physiological process depends on the fidelity of the physics-based model $\mathcal{M}$.
In addition to modeling the natural state of the physiological process, we also consider physics-based models of clinical interventions (e.g., medical devices). To model an interventional process, we consider an augmented version of the physics-based model ~$\widetilde{\mathcal{M}}:\Theta\times\mathcal{U}\to\mathcal{X}$, which maps the endogenous parameters $\theta\in\Theta$ of the patient and an exogenous intervention $u\in\mathcal{U}$ to a physiological state $\mathbf{x}\in\mathcal{X}$. The augmented model describes how the intervention $u$ alters the physiological states of a patient with parameters $\theta$—in the causal inference literature, this class of models is known as structural causal models (SCMs) (56).
To identify the model $\mathcal{M}(\theta)$ traditionally and simulate the states $\mathbf{x}$, one would need to conduct a number of invasive procedures for each patient. Here, we assume that we have access to a measurement $\mathbf{y}_{i}$ for each patient $i$, which reflects some aspects of the underlying physiological process. We assume that $\mathbf{y}$ is acquired through a non-invasive procedure that might be conducted routinely in clinical practice (e.g., an electrocardiogram). The relationship between the physiological state $\mathbf{x}$ and the non-invasive measurement $\mathbf{y}$ is given by
@@
However, we do not directly observe the state $\mathbf{x}$, but rather a measurement $\mathbf{y}$ obtained by passing $\mathbf{x}$ through the unknown forward model $\mathcal{K}$. Thus, obtaining the physical parameters from $\mathbf{y}$ entails solving the composite inverse problem given by:
@@
Solving this problem entails two challenges. First, we do not know the forward model $\mathcal{K}$. Second, we cannot directly learn the inverse map $\mathcal{F}^{-1}$ using supervised learning since this requires access to a labeled dataset $\{(\mathbf{y}_{i},\theta_{i})\}_{i}$. Obtaining such a dataset would require conducting a large number of invasive procedures, which we seek to avoid.
An alternative way to learn the inverse map $\mathcal{F}^{-1}$ is to break down the composite inverse problem in (5) into separate inverse problems, and only learn the inverse of the unknown map $\mathcal{K}^{-1}$ using examples of the form $\{(\mathbf{y}_{i},\mathbf{x}_{i})\}_{i}$ by fixing the map $\mathcal{M}$. If we can construct a mapping representative of $\mathcal{M}$ from any patient parameter set $\theta\in\Theta$ to state $\mathbf{x}\in\mathcal{X}$, then as we can write $\mathcal{K}^{-1}$ as follows:
$\mathcal{K}^{-1}:\mathcal{Y}\to_{\mathcal{F}^{-1}}\Theta\to_{\mathcal{M}}\mathcal{X}$ (6)
we can learn $\mathcal{F}^{-1}$ using training samples which only train $\mathcal{K}^{-1}$.
As we will ultimately use supervised methods and gradient descent to learn Eq. (6), we need a method for fixing $\mathcal{M}$ that allows for mapping any arbitrary parameter set $\theta$ to corresponding state $\mathbf{x}$. There are several methods in the literature for solving the forward dynamics of $\mathcal{M}$ that we might think of to construct this mapping. These include physics-informed neural networks (PINNs) (59), which train a network to approximate the solution $\mathbf{x}_{i}$ to $f_{\theta_{i}}(\mathbf{x}_{i})=0$ thereby solving $\mathcal{M}(\theta_{i})$ for a fixed patient parameter set $\theta_{i}$. Traditional PINNs learn the function $\mathbf{x}_{i}$ by incorporating the differential equation $f_{\theta_{i}}=0$ as a soft-constraint in the network’s training loss alongside a data loss. In the setting of a ordinary differential equations, Neural ODEs (10) train a network to learn the differential equation operator $f_{\theta_{i}}$ on the states $\mathbf{x}_{i}$ to solve for $\mathbf{x}_{i}=\mathcal{M}(\theta_{i})$ using differentiable ODE solvers. These methods can learn $\mathcal{M}(\theta_{i})$ for a fixed patient parameter set $\theta_{i}$ and can be compared to the use of a numerical differential equation solvers.
These existent approaches for solving $\mathcal{M}$, however, only give the states $\mathbf{x}$ for a single patient at a time. To learn $\mathcal{K}^{-1}$ by training a network through the map $\mathcal{M}$, we need a method that approximates the map as otherwise, we would need to train infinitely many separate PINNs or Neural ODEs to cover the entirety of a plausible parameter range potentially traversed in supervised learning of (6).
Non-Invasive Data Acquisition. While we do not have direct access to the patient physiological state $\mathbf{x}_{i}$ to train a model to learn $\mathcal{K}^{-1}$, the noninvasive observation $\mathbf{y}_{i}$ typically measures physiological quantities that can also be derived from the true state $\mathbf{x}_{i}$. We assume that $\bar{\mathbf{x}}$ is a physiological variable that can be derived from both $\mathbf{x}$ and $\mathbf{y}$, i.e.,
@@
Given the setup above, our digital twin modeling problem can be formulated as follows. Given a dataset of noninvasive measurements for $n$ patients, $\mathcal{D}=\{\mathbf{y}_{i}\}_{i=1}^{n}$, and a physiological variable of interest $\bar{\mathbf{x}}_{i}=m(\mathbf{x}_{i})=g(\mathbf{y}_{i})$, our goal is to train a model that can identify the underlying physics-based twin $\mathcal{M}(\theta_{n+1})$ for a new patient $n+1$ based solely on their corresponding noninvasive measurement $\mathbf{y}_{n+1}$. We seek to do this by fixing $\mathcal{M}$ and learning $\mathcal{K}^{-1}$ from the data $\mathcal{D}=\{(\mathbf{y}_{i},\bar{\mathbf{x}}_{i})\}_{i=1}^{n}$.
@@ 2.2 Physics-Informed Self-Supervised Learning
The motivation for training the model $\widehat{f}$ is usually to automate the collection of physiological variables from medical images in a clinical workflow (67). These models follow a fully data-driven approach and do not incorporate our knowledge of how the physiological processes being measured function on the biological or physical levels. One could think of the supervised model $\widehat{f}$ as a data-driven model of the function $g(.)$ in (7).
@@
$\widehat{\phi}_{\mathcal{M}}=\arg\min_{\phi}\left\{\frac{1}{\tilde{n}}\sum_{j=1}^{\tilde{n}}\ell(\phi(\tilde{\theta}_{j}),\tilde{\mathbf{x}}_{j})\right\}.$ (11)
@@
Step 2: Physics-Guided Finetuning. Given the observed dataset $\mathcal{D}=\{(\mathbf{y}_{i},\bar{\mathbf{x}}_{i})\}_{i}$, we train another model to predict the physical parameters $\theta$ from the observed measurements $\mathbf{y}$ using the loss function
@@
Interpretation of Physics-Informed SSL. Our proposed algorithm structurally resembles the SSL paradigm, where, akin to SSL, we decompose our model into a “backbone” architecture and a task-specific “head” (46). However, there are fundamental conceptual differences between standard SSL approaches and ours. In standard SSL, the backbone is a high-capacity model pretrained on a pretext task using unlabeled data to derive a general-purpose representation transferable to many downstream tasks, while the low-capacity head is finetuned for the specific task. In our physics-informed approach, SSL serves to constrain rather than enhance flexibility. We first pretrain the low-capacity head to distill the laws of physics, and then finetune the high-capacity backbone model to learn a mapping that aligns with the frozen head to produce the observed physiological variables $\mathbf{x}$ while respecting the physical laws. In doing so, we force the backbone model to learn a mapping from the measurement $\mathbf{y}$ to the physical parameters $\theta$.
The concept of distilling a physics-based dynamic model into a data-driven one draws from engineering methodologies that develop surrogate models to simplify computational complexity using simulated data (e.g., (57)). This data-driven approach mirrors other supervised learning generative strategies for solving inverse problems or system identification with known dynamics in control theory (50; 77). Importantly, physics-informed SSL scales efficiently with increasingly complex and higher fidelity physical models, as its computational requirements involve one offline process (Eq. (10) and (11)) once the synthetic data is acquired.
Figure 2: Illustration of Med-Real2Sim digital twins for cardiovascular hemodynamics. (a) Pictorial depiction of the two-step physics-informed SSL algorithm proposed in Section 2.2. (b) Five state lumped-parameter electric circuit model of cardiac hemodynamics from (65). Here $\mathbf{x}(t)=[x_{1}(t),x_{2}(t),x_{3}(t),x_{4}(t),x_{5}(t)]=$ $[P_{LV}(t),P_{LA}(t),P_{A}(t),P_{Ao}(t),Q(t)]$ describes the voltages $x_{1},x_{2},x_{3},x_{4}$ or pressures in the left-ventricle, left atrium, arteries, and aorta, respectively, and total flow $x_{5}$. The LVAD is modeled through an electric circuit connected to the digital twin via a switch. An LVAD intervention is applied if the switch is closed.